A dielectric object is placed in an electric field. The object becomes polarised and 

a large number of atomic/molecular dipoles in the object align in the direction of 

the applied electric field. Derive an expression for the electric potential due to this 

polarised dielectric at a point outside the dielectric.


1
Expert's answer
2022-02-23T08:33:10-0500

"\Delta\varphi=\frac 1{4\pi\varepsilon_0}\frac{q^--q^+}{r_x^2+r_y^2}."